# Accurate estimation of feature importance faithfulness for tree models

### v1

@@ Abstract
In this paper, we consider a perturbation-based metric of predictive faithfulness of feature rankings (or attributions) that we call PGI squared. When applied to decision tree-based regression models, the metric can be computed accurately and efficiently for arbitrary independent feature perturbation distributions. In particular, the computation does not involve Monte Carlo sampling that has been typically used for computing similar metrics and which is inherently prone to inaccuracies.
Moreover, we propose a method of ranking features by their importance for the tree model’s predictions based on PGI squared. Our experiments indicate that in some respects, the method may identify the globally important features better than the state-of-the-art SHAP explainer.
1 Introduction
One of the key challenges in deploying modern machine learning models in such areas as medical diagnosis lies in the ability to indicate why a certain prediction has been made. Such an indication may be of critical importance when a human decides whether the prediction can be relied on. This is one of the reasons various aspects of explainability of machine learning models have been the subject of extensive research lately (see, e.g., [3]).
For some basic types of models (e.g., single decision trees), the rationale behind a prediction is easy to understand by a human. However, predictions of more complex models (that offer much better accuracy, e.g., based on neural networks or decision tree ensembles) are also much more difficult to interpret. Accurate and concise explanations understandable to humans might not always exist. In such cases, it is still beneficial to have methods giving a flavor of what factors might have influenced the prediction the most.
@@ 1 Introduction
Many local attribution methods have been proposed so far, e.g., [12, 14, 15]. They are often completely model-agnostic, that is, they can be defined for and used with any model which is only accessed in a black-box way for computing the attribution. In some of these cases, focusing on a particular model’s architecture can lead to much more efficient and accurate attribution computation algorithms (e.g., [10] for tree ensemble models).
@@
With the large body of different attribution methods, it is not clear which one should one use, especially since there is not a single objective measure of their reliability. To help deal with this issue, several notions of explanation quality have been proposed in the literature, such as faithfulness (fidelity) [11, 17], stability [2], or fairness [4, 8]. [1] developed an open-source benchmark OpenXAI automatically computing variants of these measures for a number of proposed attribution methods.
Out of these notions, our focus in this paper is on faithfulness.
@@
Roughly speaking, an attribution method deserves to be called faithful if the features deemed important by the attribution are truly important for the decision-making process of the model. The perturbation-based methods constitute one popular class of approaches to measuring faithfulness. Intuitively, perturbing features deemed impactful should generally lead to a significant change in the prediction. On the contrary, manipulating unimportant features should not make a big difference.
@@
One concrete cleanly-defined perturbation-based faithfulness metric for regression problems is the prediction gap on important feature perturbation, PGI in short. PGI is a faithfulness measure of choice in the OpenXAI benchmark [1]. Having fixed some subset of important features $S$ derived from the obtained attribution, the prediction gap wrt. $S$ is defined as the expected absolute change in the prediction when an independent Gaussian noise is applied to each coordinate $x_{j}$, $j\in S$. PGI can be defined as either the prediction gap wrt. a fixed set of important features (e.g., $k$ top-scoring features for some fixed $k$ [8]), or an average prediction gap over many important feature sets (as in [1]). In particular, the PGI metric depends only on the ordering of features by importance induced by the attribution. Other perturbation-based methods include e.g., PGU [1], comprehensiveness, sufficiency [7], remove-and-retrain [9], deletion [13].
@@
In practice, PGI and other methods using random perturbations are computed using the Monte Carlo method which is inherently prone to inaccuracies, unless a very large number of samples is used.
@@
It is interesting to ask whether PGI or similar random perturbation-based faithfulness matrices can be computed more accurately, say, with the approximation error caused merely by the floating-point arithmetic’s rounding errors.
1.1 Our contribution
1.1.1 Exact computation
@@
Unfortunately, computing the prediction gap, as defined in [1] (wrt. a fixed set $S$ of perturbed features, see Section 2 for a formal definition), does not look very tractable given the rich structure that tree ensemble models offer. This is because therein, the expectation is taken over the absolute value of $f(x)-f(x^{\prime})$, where $x^{\prime}$ is $x$ with important features perturbed. The absolute value is not very mathematically convenient to work with when combined with expectations. On the other hand, dropping the absolute value here would not make much sense: a small random perturbation (positive or negative) of a feature could in principle lead to a large positive or negative prediction difference. These, in turn, could cancel out in expectation, making the metric close to zero even though the feature is visibly of high importance.
@@
To deal with the tractability problem of measuring PGI on tree models, we propose a similar metric, $\text{PGI}^{2}$, where instead of taking the absolute value, we compute the expectation of the square prediction gap $(f(x)-f(x^{\prime}))^{2}$. This eliminates the cancellation problem outlined above. Of course, introducing the square can alter the evaluation of the feature’s importance wrt. PGI as defined previously. However, intuitively, it amplifies the big prediction differences, and at the same time marginalizes the small, which is a generally good property if we seek concise explanations.
@@
We prove that for tree ensemble models with $n$ nodes in total, the squared prediction gap can be computed in $O(n^{2})$ time exactly for any fixed choice of subset of perturbed important features, assuming the cumulative distribution function of feature-wise random perturbations can be evaluated in constant time. That is, the obtained running time bound does not require that the random perturbations are Gaussian or their distributions are equal across features.
@@
The OpenXAI benchmark calculates PGI by taking the average expected prediction gap obtained when perturbing $1,2,\ldots,d$ most important features (where $d$ denotes the total number of features; see  [1, Appendix A]). In such a case, $d$ different expectations of absolute prediction gap are estimated, each of them via Monte Carlo sampling. In our case, $\text{PGI}^{2}$ is defined as the average expected squared prediction gap when perturbing $1,2,\ldots,d$ most important features. Using our prediction gap algorithm, $\text{PGI}^{2}$ is computed exactly (assuming infinite precision arithmetic; in reality, up to the precision error incurred by the usage of floating-point arithmetic) in $O(n^{2}d)$ time.
Experiments.
We accompany our theoretical development with an experimental comparison of the results obtained using the exact algorithm and Monte Carlo sampling.
The experiments confirm that Monte Carlo simulation gives noticeably less accurate results within the allocated time (iteration) budget comparable to the running time of our algorithm. Specifically, in our experiments, the normalized mean absolute error (NMAE) is typically between $0.02$ and $0.3$. The error measured this way generally decreases when increasing the size of the model which follows from the quadratic complexity of our algorithm. It is also decreasing in the magnitude of perturbations used: sampling is generally better at estimating larger quantities, whereas tiny perturbations lead to tiny prediction gaps.
1.1.2 $\text{PG}^{2}$-based greedy feature ordering.
@@
We consider constructing the feature ranking using a greedy $\text{PG}^{2}$ heuristic: the $i$-th most important feature $f$ is chosen so that it optimizes the prediction gap together with the already chosen $i-1$ most important features plus $f$.
We next compare the greedy $\text{PG}^{2}$ ranking with the ranking produced using the popular SHAP feature attribution method for tree ensembles [10]. First of all, our experiments confirm that the greedy $\text{PG}^{2}$ ranking yields, on average, better $\text{PGI}^{2}$ scores than SHAP.
While that such a property holds is non-obvious, it is perhaps not very surprising either. This is why we also compare our ranking method with SHAP in terms of average conciseness. Specifically, we consider three natural metrics of averaged feature importance across the test data set, each of them based solely on the feature ordering and clearly promoting the most important features. For each of these metrics, we compare the entropy of the obtained averaged importances. Intuitively, a smaller entropy indicates that only few features are identified as critically important to the model.
Our experiments show that the greedy $\text{PG}^{2}$ ranking used with perturbations of sufficiently large standard deviation can lead to global importance scores with smaller entropy compared to SHAP.
2 Preliminaries
@@
Prediction gap on important feature perturbation.
Let a ranking $\pi\in[d]\to[d]$ be a permutation of features giving their importance from highest to lowest. Denote by $\pi[1..\ell]$ the set $\{\pi(1),\ldots,\pi(\ell)\}$, that is, the $\ell$ top ranking features in $\pi$. For any $x\in\mathbb{R}^{d}$ and $S\subseteq[d]$, let $\text{perturb}(x,S)$ be some distribution on feature vectors $x^{\prime}\in\mathbb{R}^{d}$ satisfying $x^{\prime}_{i}=x_{i}$ for all $i\notin S$. AKS+ [1] define:
$\mathrm{PGI}(x,\pi,k):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[|f(x^{\prime})-f(x)|\right].$
Based on that, they define an averaged metric:
$\mathrm{PGI}(x,\pi):=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}(x,\pi,k).$
We analogously define the $\mathrm{PGI}^{2}$ metric:
$\mathrm{PGI}^{2}(x,\pi,k)$ $:=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[\left(f(x^{\prime})-f(x)\right)^{2}\right],$
$\mathrm{PGI}^{2}(x,\pi)$ $:=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}^{2}(x,\pi,k).$
3 Computing $\mathrm{PGI}^{2}$ for tree ensemble models
Consider distributions $\text{perturb}(x,S)$ such that ${x^{\prime}\sim\text{perturb}(x,S)}$ satisfies:
@@ 3 Computing $\mathrm{PGI}^{2}$ for tree ensemble models
    for every $i\in S$, $x^{\prime}_{i}=x_{i}+\delta_{i}$, where $\delta_{i}\sim\mathcal{D}_{i}$, and
@@
Moreover, we assume that for all $i\in S$, the c.d.f. $F_{\mathcal{D}_{i}}$ of $\mathcal{D}_{i}$ can be evaluated in $O(1)$ time.
In the following, we show how to compute
$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right]$ (1)
under the assumptions made in the case when $f$ is a tree ensemble model $(\mathcal{T})_{i=1}^{m}$. Note that the metric $\mathrm{PGI}^{2}(x,\pi)$ for a given ranking $\pi$ can be computed by running the algorithm  $d$ times for different subsets $S$ of the form $\pi[1..k]$.
Put $c=f(x)$. Let $x^{\prime}\sim\text{perturb}(x,S)$. Recall that ${f(x^{\prime})=\sum_{i=1}^{m}f_{\mathcal{T}_{i}}(x^{\prime})}$. For every leaf node $v$ of some $\mathcal{T}_{i}$, consider a random indicator variable $X_{v}$ such that $X_{v}=1$ iff $\mathcal{T}_{i}$ evaluates to the value $y_{v}$ of leaf $v$ given $x^{\prime}$ as input. In particular, for any $i=1,\ldots,m$, there is exactly one $v\in\mathcal{L}(\mathcal{T}_{i})$ such that $X_{v}=1$. Set $\mathcal{L}=\bigcup_{i=1}^{m}\mathcal{L}(\mathcal{T}_{i})$. Then:
@@
It follows that $X_{u}=1\land X_{v}=1$ iff for all $q\in Q_{u}\cup Q_{v}$, $x^{\prime}_{q}\in I_{u,q}\cap I_{v,q}$. Hence, if we denote by $l_{u,v,q},r_{u,v,q}$ the respective endpoints of $I_{u,q}\cap I_{v,q}$, by the independence of all perturbations $\delta_{j}$, we obtain that $\Pr[X_{u}=1\land X_{v}=1]$ equals $\Pi(u,v)$, where
$\Pi(u,v)=\left(\prod_{q\in(Q_{u}\cup Q_{v})\setminus S}[x_{q}\in I_{u,q}\cap I_{v,q}]\right)\cdot\left(\prod_{q\in(Q_{u}\cup Q_{v})\cap S}\left(F_{\mathcal{D}_{q}}(r_{u,v,q}-x_{q})-F_{\mathcal{D}_{q}}(l_{u,v,q}-x_{q})\right)\right).$ (2)
@@
Let $f$ be a tree ensemble model whose trees have $n$ nodes in total. Let $x\in\mathbb{R}^{d}$. Then for any $S\subseteq[d]$, $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time. In particular, given a feature ranking $\pi$, for any $k\in[d]$, $\mathrm{PGI}^{2}(x,\pi,k)$ can be computed in $O(n^{2})$ time, whereas $\mathrm{PGI}^{2}(x,\pi)$ can be computed in $O(n^{2}d)$ time.
4 Experiments
4.1 Datasets
Following the previous work (e.g., [1, 8]), in the experiments we assume that the (real) perturbations used for computing the prediction gap (either exactly, or approximately) come from the same normal distribution $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$. This is why we need to assume that all the features have numerical values.
Using the same perturbation distribution for all the features requires the features be standardised by subtracting the mean and dividing by the standard deviation. This ensures that if a feature is perturbed by a noise variable drawn from $\mathcal{D}$, the perturbation will have a similar effect. More specifically, it prevents a perturbation from $\mathcal{D}$ being relatively significant for one feature and insignificant for another.
In our experiments we used two datasets:
@@ 4.1 Datasets
    Red Wine Quality Dataset [5] The dataset contains 11 features related to physico-chemical attributes of wine, all numerical and continuous. The task is to predict the score of a wine, which is an integer between 1 and 10. The task can be considered as both a regression and a classification task. The dataset contains 1 59915991\ 5991 599 examples.
@@
    California Housing [16] The dataset contains information from the 1990 Californian census on some characteristics of a given Californian county and the median house value in that county. There are 8 numerical characteristics, such as length or number of bedrooms, and one categorical characteristic - proximity to the ocean. For the reasons outlined before, we decided to drop this feature and use a modified dataset. The task is to predict the median value of the house. The dataset contains 20 6402064020\ 64020 640 examples.
4.2 Models
All tree ensembles are implemented in the XGBoost library [6]. For each of the two datasets, three models of the following types were trained:
  • Single tree – a single decision tree with max_depth set to 4.
  • Bigger – gradient boosted trees with max_depth set to 4 and number of trees limited to 40.
In each case, the data set was split 80:20 into training and test sets. In addition, for each model and its constraint, a grid search was performed over selected hyperparameters. The details of the training process can be found in Appendix B.
4.3 Comparing squared prediction gap algorithms
In this section we compare our algorithm computing the squared prediction gap (as defined in Equation (1)) for tree ensembles described in Section 3 and the Monte Carlo sampling-based method. Specifically, the latter algorithm, given $x$ and the set $S\subseteq[d]$ or perturbed features, evaluates $f(x^{\prime})$ for $i$ randomly sampled $x^{\prime}\sim\text{perturb}(x,S)$ (see Section 3) and records the average value of $(f(x^{\prime})-f(x))^{2}$. We will call $i$ the number of iterations in the following.
The subject of the comparison was the accuracy of the results produced by the sampling algorithm. Since the accuracy of the sampling algorithm increases with the number of iterations $i$, to make the comparison fair, the number $i$ was chosen so that the wall time of both algorithms was close. Both algorithms were implemented in Python and all the arithmetic operations in the algorithms were carried out using the numpy.float32 type.

### v3

# AAAI Press Formatting Instructions for Authors Using — A Guide

@@ Abstract
In this paper, we consider a perturbation-based metric of predictive faithfulness of feature rankings (or attributions) that we call PGI squared. When applied to decision tree-based regression models, the metric can be computed exactly and efficiently for arbitrary independent feature perturbation distributions. In particular, the computation does not involve Monte Carlo sampling that has been typically used for computing similar metrics and which is inherently prone to inaccuracies.
As a second contribution, we proposed a procedure for constructing feature ranking based on PGI squared. Our results indicate the proposed ranking method is comparable to the widely recognized SHAP explainer, offering a viable alternative for assessing feature importance in tree-based models.
Codehttps://github.com/rapicki/prediction-gap
Extended versionhttps://arxiv.org/abs/2404.03426
Introduction
One of the key challenges in deploying modern machine learning models in such areas as medical diagnosis lies in the ability to indicate why a certain prediction has been made. Such an indication may be of critical importance when a human decides whether the prediction can be relied on. This is one of the reasons various aspects of explainability of machine learning models have been the subject of extensive research lately (see, e.g., (Burkart and Huber 2021)).
For some basic types of models (e.g., single decision trees, the rationale behind a prediction is easy to understand by a human. However, predictions of more complex models (that offer much better accuracy, e.g., based on neural networks or decision tree ensembles) are also much more difficult to interpret. Accurate and concise explanations understandable to humans might not always exist. In such cases, it is still beneficial to have methods giving a flavor of what factors might have influenced the prediction the most.
@@ Introduction
Many local attribution methods have been proposed so far, e.g., (Lundberg and Lee 2017; Plumb, Molitor, and Talwalkar 2018; Ribeiro, Singh, and Guestrin 2016). They are often completely model-agnostic, that is, they can be defined for and used with any model which is only accessed in a black-box way for computing the attribution. In some of these cases, focusing on a particular model’s architecture can lead to much more efficient and accurate attribution computation algorithms (e.g., (Lundberg et al. 2020) for tree ensemble models).
@@
With the large body of different attribution methods, it is not clear which one should one use, especially since there is not a single objective measure of their reliability. To help deal with this issue, several notions of explanation quality have been proposed in the literature, such as faithfulness (fidelity) (Liu et al. 2021; Yeh et al. 2019), stability (Alvarez-Melis and Jaakkola 2018), or fairness (Balagopalan et al. 2022; Dai et al. 2022). Agarwal et al. (2022) developed an open-source benchmark OpenXAI automatically computing variants of these measures for a number of proposed attribution methods.
@@
Out of these notions, our focus in this paper is on faithfulness. Roughly speaking, an attribution method deserves to be called faithful if the features deemed important by the attribution are truly important for the decision-making process of the model. The perturbation-based methods constitute one popular class of approaches to measuring faithfulness. Intuitively, perturbing features deemed impactful should generally lead to a significant change in the prediction. On the contrary, manipulating unimportant features should not make a big difference.
@@
One concrete cleanly-defined perturbation-based faithfulness metric for regression problems is the prediction gap on important feature perturbation, PGI in short. PGI is a faithfulness measure of choice in the OpenXAI benchmark (Agarwal et al. 2022). Having fixed some subset of important features $S$ derived from the obtained attribution, the prediction gap (PG) of $x$ wrt. $S$ is defined as:
$\mathrm{PG}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[|f(x^{\prime})-f(x)|\right].$ (1)
In (Agarwal et al. 2022), the concrete perturbation method $\text{perturb}(x,S)$ is adding an independent Gaussian noise from $\mathcal{N}(0,\sigma^{2})$ to each coordinate $x_{j}$, $j\in S$. $\mathrm{PGI}(x)$ can be defined as either the prediction gap wrt. a fixed set of important features (e.g., $k$ top-scoring features for some fixed $k$ (Dai et al. 2022)), or an average prediction gap over many important feature sets (as in (Agarwal et al. 2022)). In particular, the PGI metric depends only on the ordering of features by importance induced by the attribution. Other perturbation-based methods include e.g., PGU (Agarwal et al. 2022), comprehensiveness, sufficiency (DeYoung et al. 2020), remove-and-retrain (Hooker et al. 2019), deletion (Petsiuk, Das, and Saenko 2018).
@@
In practice, PGI and other methods based on random perturbations are computed using the Monte Carlo method which is inherently prone to inaccuracies, unless a large number of samples is used.
@@
It is interesting to ask whether PGI or similar random perturbation-based faithfulness metrics can be computed exactly, that is, via a closed-form calculation, so that any approximation error is caused merely by the floating-point arithmetic’s rounding errors.
Our Contribution
Exact Computation
@@
Unfortunately, computing $\mathrm{PG}(x,S)$ (wrt. a fixed set $S$ of perturbed features, as in (Agarwal et al. 2022), see (1)) does not look very tractable even given the rich structure that tree ensemble models offer. This is because therein, the expectation is taken over the absolute value of $f(x)-f(x^{\prime})$, where $x^{\prime}$ is $x$ with important features perturbed. The absolute value is not very mathematically convenient to work with when taking expectations. On the other hand, dropping the absolute value here would not make much sense: a small random perturbation (positive or negative) of a feature could in principle lead to a large positive or negative prediction difference. These, in turn, could cancel out in expectation, making the metric close to zero even though the feature is visibly of high importance.
Theoretical contribution: To deal with the tractability problem of measuring PGI on tree models, we propose a similar metric, $\mathrm{PGI}^{2}$, obtained from a slightly different squared prediction gap:
$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right].$ (2)
@@
This eliminates the cancellation problem outlined above. Of course, introducing the square can alter the evaluation of the feature’s importance wrt. PGI as defined previously. However, intuitively, it amplifies the big prediction differences, and at the same time marginalizes the small, which is a generally good property if we seek concise explanations.
@@
We prove that for tree ensemble models with $n$ nodes in total, the squared prediction gap $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time exactly for any fixed choice of a subset $S$ of perturbed important features, assuming the cumulative distribution function of feature-wise random perturbations can be evaluated in constant time. That is, the obtained running time bound does not require that the random perturbations are Gaussian or their distributions are equal across features.
We note that whereas our $\mathrm{PG}^{2}$ algorithm is quadratic, from the theoretical point of view, the important thing is that on tree models, a prediction gap-style quantity computation can be carried out exactly in polynomial time after all.
@@
Given a permutation $\pi$ ranking the $d$ features from the most to least important, the OpenXAI benchmark calculates $\mathrm{PGI}$ by taking the average prediction gap obtained when perturbing $k=1,2,\ldots,d$ most important features $\pi[1..k]=\{\pi(1),\ldots,\pi(k)\}$ (see (Agarwal et al. 2022, Appendix A)). In such a case, $d$ different expectations of absolute prediction gap are estimated, each via Monte Carlo sampling. In our case, $\mathrm{PGI}^{2}(x)$ is analogously defined as:
$\mathrm{PGI}^{2}(x,\pi):=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PG}^{2}(x,\pi[1..k]).$ (3)
Using our prediction gap algorithm, $\mathrm{PGI}^{2}$ is computed exactly (assuming infinite precision arithmetic; in reality, up to the precision error incurred by the usage of floating-point arithmetic) in $O(n^{2}d)$ time.
Experiments: In our experiments, we evaluated different methods for calculating the Prediction Gap. Our study encompassed the exact algorithm as well as two sampling techniques integration techniques: Monte Carlo (MC) and Quasi-Monte Carlo (QMC). With the increase in iteration count, the outputs of MC and QMC visibly converged to the output of our exact algorithm; this confirms good numerical stability of our approach.
When allocating an equivalent time budget across all methods, our findings revealed an advantage over MC and QMC, with QMC demonstrating a slight edge. Specifically, the Normalized Mean Absolute Error (NMAE) for MC was 0.13 for single models and decreased to 0.01 for bigger models. In comparison, QMC exhibited NMAEs of approximately 0.05 for single models and 0.002 for bigger ones. These results underscore the efficacy of sampling methods in the context of computing the PG, particularly for more sophisticated model structures.
$\text{PG}^{2}$-Based Greedy Feature Ordering.
@@
We consider constructing the feature ranking using a greedy $\mathrm{PG}^{2}$ heuristic: the $i$-th most important feature $\lambda$ is chosen so that it optimizes the squared prediction gap together with the already chosen $i-1$ most important features plus $\lambda$. We next compare the greedy $\mathrm{PG}^{2}$ ranking with the ranking produced using the popular SHAP feature attribution method for tree ensembles (Lundberg et al. 2020). We observe that the most important features identified by the two methods generally differ, and the deviation is clearer if the applied perturbations are smaller.
As far as faithfulness is considered, our experiments confirm that the greedy $\mathrm{PG}^{2}$ ranking yields, on average, better $\mathrm{PGI}^{2}$ scores than SHAP. While that such a property holds is non-obvious, it is perhaps not very surprising either. This is why we also compare the $\mathrm{PG}^{2}$-based ranking method with SHAP in terms of two different faithfulness metrics: feature randomization and remove-and-retrain (Hooker et al. 2019).
In the former case and some of our datasets, the $\mathrm{PG}^{2}$-based rankings (used with Gaussian perturbations for small standard deviations) performed better than SHAP, whereas in the others, the results were comparable. Wrt. the latter metric, SHAP achieved better performance.
Preliminaries
@@
Computing $\mathrm{PG}^{2}$ for Tree Ensemble Models
Let $S\subseteq[d]$ be a subset of important features and $x\in\mathbb{R}^{d}$. In this section, we show our polynomial exact algorithm computing $\mathrm{PG}^{2}(x,S)$, as defined in (2).
We consider distributions $\text{perturb}(x,S)$ such that ${x^{\prime}\sim\text{perturb}(x,S)}$ satisfies:
@@
    for every perturbed feature $i\in S$, $x^{\prime}_{i}=x_{i}+\delta_{i}$, where $\delta_{i}\sim\mathcal{D}_{i}$, and distribution $\mathcal{D}_{i}$ is fixed, and the c.d.f. $F_{\mathcal{D}_{i}}$ of $\mathcal{D}_{i}$ can be evaluated in $O(1)$ time.
@@
The main goal of this section is to show how to compute
$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right]$
under the assumptions made in the case when $f$ is a tree ensemble model $(\mathcal{T})_{i=1}^{m}$. Recall from Equation (3) that the metric $\mathrm{PGI}^{2}(x,\pi)$ for a given ranking $\pi$ can be computed by running the algorithm $d$ times for $d$ different subsets $S$ of the form $\pi[1..k]$.
Put $c:=f(x)$. Let $x^{\prime}\sim\text{perturb}(x,S)$. Recall that ${f(x^{\prime})=\sum_{i=1}^{m}f_{\mathcal{T}_{i}}(x^{\prime})}$. For every leaf node $v$ of some $\mathcal{T}_{i}$, consider a random indicator variable $X_{v}$ such that $X_{v}=1$ iff $\mathcal{T}_{i}$ evaluates to the value $y_{v}$ of leaf $v$ given $x^{\prime}$ as input. In particular, for any $i=1,\ldots,m$, there is exactly one $v\in\mathcal{L}(\mathcal{T}_{i})$ such that $X_{v}=1$. Set $\mathcal{L}=\bigcup_{i=1}^{m}\mathcal{L}(\mathcal{T}_{i})$. Then:
@@
It follows that the event $X_{u}=1\land X_{v}=1$ occurs iff for all $q\in Q_{u}\cup Q_{v}$, $x^{\prime}_{q}\in I_{u,q}\cap I_{v,q}$. Hence, if we denote by $l_{u,v,q},r_{u,v,q}$ the respective endpoints of $I_{u,q}\cap I_{v,q}$, by the independence of all perturbations $\delta_{j}$, we obtain that $\Pr[X_{u}=1\land X_{v}=1]$ equals $\Pi(u,v)$, where
$\begin{split}\Pi(u,v)=\left(\prod\limits_{q\in(Q_{u}\cup Q_{v})\setminus S}[x_{q}\in I_{u,q}\cap I_{v,q}]\right)\cdot\\ \prod\limits_{q\in(Q_{u}\cup Q_{v})\cap S}\left(F_{\mathcal{D}_{q}}(r_{u,v,q}-x_{q})-F_{\mathcal{D}_{q}}(l_{u,v,q}-x_{q})\right).\end{split}$ (4)
@@
This idea is the realized in Algorithm 1, where, while the two nested pre-order traversals over $\mathcal{T}$ proceed, the endpoints of intervals $I_{\cdot,\cdot}$ are maintained using the arrays $l,r$, and the factors of (4) are maintained using the array factor.
Algorithm 1 Computing $\Pi(u,v)$ for all leaf pairs $u,v$ given model $\mathcal{T}$, a feature vector $x$, and important features $S\subseteq[d]$
$\forall_{q\in[d]}\quad l[q]=\infty,r[q]=\infty,\text{factor}[q]=1$
2:procedure Compute-Prob($v,\text{prod}=1,u=\textbf{null}$)
3:     if $v$ is non-root $\triangleright$ update the maintained arrays
4:         $p:=$ parent of $v$ in $\mathcal{T}$
5:         $q:=q_{p}$
6:         $(l^{\prime},r^{\prime},f^{\prime}):=(l[q],r[q],\text{factor}[q])$
7:         if $v=a_{p}$ $\triangleright$ $v$ is the left child of its parent
8:              $r[q]:=\min(r[q],t_{p})$
9:         else
10:              $l[q]:=\max(l[q],t_{p})$
11:         end if
12:         if $q\in S$ $\triangleright$ See Equation (4)
13:              $\text{factor}[q]:=F_{\mathcal{D}_{q}}(r[q]-x_{q})-F_{\mathcal{D}_{q}}(l[q]-x_{q})$
14:         else
15:              $\text{factor}[q]:=[x_{q}\geq l[q]\land x_{q}\leq r[q]]$
16:         end if
17:         $\text{prod}:=(\text{prod}/f^{\prime})\cdot\text{factor}[q]$
18:     end if
19:     if $v$ is a leaf then
20:         if $u=\textbf{null}$ then
21:              for $T\in\mathcal{T}$ $\triangleright$ inner loop through trees
22:                  Compute-Prob($\text{root}(T),\text{prod},v$)
23:              end for
24:         $\triangleright$ both leaves $u$ and $v$ are fixed
25:              $\Pi(u,v):=\text{prod}$
26:         end if
27:     $\triangleright$ descend
28:         Compute-Prob($a_{v},\text{prod},u$)
29:         Compute-Prob($b_{v},\text{prod},u$)
30:     end if
31:     if $v$ is non-root $\triangleright$ revert changes to $l,r,\text{factor}$
32:         $(l[q],r[q],\text{factor}[q]):=(l^{\prime},r^{\prime},f^{\prime})$
33:     end if
34:end procedure
35:
36:for $T\in\mathcal{T}$ $\triangleright$ outer loop through trees
37:     Compute-Prob($\text{root}(T)$)
38:end for
@@
Let $f$ be a tree ensemble model whose trees have $n$ nodes in total. Let $x\in\mathbb{R}^{d}$. Then for any $S\subseteq[d]$, $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time.
Experiments: Exact vs Monte Carlo Sampling
In this section, we compare our algorithm computing the squared prediction gap (as defined in Equation (2)) for tree ensembles as described earlier, the regular Monte Carlo sampling-based method (MC) and Quasi Monte Carlo sampling method (QMC) [^2] The details of quasi-random generators used in our Quasi Monte Carlo implementation can be found in the Appendix.. Specifically, the Monte Carlo methods, given $x$ and the set $S\subseteq[d]$ or perturbed features, evaluates $f(x^{\prime})$ for $i$ randomly sampled $x^{\prime}\sim\text{perturb}(x,S)$ and records the average value of $(f(x^{\prime})-f(x))^{2}$ over the samples. We will call $i$ the number of iterations in the following.
Recall that the accuracy of Monte Carlo methods is supposed to increase with the number of iterations $i$. The subject of the comparison was hence the accuracy of the results produced by the sampling algorithms as a function of the number of iterations with the goal of deciding which method (exact, MC, or QMC) is preferable and under what conditions.
Specifically, having fixed a model $f$, we looked at the average difference between the values $\mathrm{PG}^{2}(x,S)$ computed using our closed-form algorithm and the two considered sampling methods. Using the output of our algorithm as the “ground truth” requires the algorithm to be numerically stable. So an indirect goal of the comparison was to establish whether our algorithm has good numerical properties.
Experimental Setup
Datasets
Following the previous work (e.g., (Agarwal et al. 2022; Dai et al. 2022)), in each experiment we assume that the (real) perturbations used for computing the prediction gap (either exactly, or approximately) come from the same normal distribution $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$. This is why we need to assume that all the features have numerical values.
Using the same perturbation distribution for all the features requires the features be standardized by subtracting the mean and dividing by the standard deviation. This ensures that if a feature is perturbed by a noise variable drawn from $\mathcal{D}$, the perturbation will have a similar effect. More specifically, it prevents a perturbation from $\mathcal{D}$ being relatively significant for one feature and insignificant for another.
In our experiments, we used the following datasets:
@@
    Red Wine Quality (Cortez et al. 2009). The dataset contains 11 features wine, all numerical and continuous. The task is to predict the score of a wine, which is an integer between 1 and 10. We considered it as a regression task. The dataset contains 1 59915991\ 5991 599 examples and has 11 features.
@@
    California Housing (Torgo 2023) The dataset contains information from the 1990 Californian census. There are 8 numerical characteristics, and one categorical - proximity to the ocean. For the reasons outlined before, we decided to drop this feature and use a modified dataset. The task is to predict the median value of the house. The dataset contains 20 6402064020\ 64020 640 examples and has 8 features.
  3. 3.
    Parkinson Telemonitoring Data (Tsanas and Little 2009) The dataset contains 5 87558755\ 8755 875 voice measurements from Parkinson’s disease patients, collected at home. It includes 17 numerical features, after dropping 3 categorical columns(ID, age) with a task to predict UPDRS motor and total scores.
Models
All tree ensembles are implemented in the XGBoost library (Chen and Guestrin 2016). For each of the datasets, models of the following two types were trained:
  • Single tree – one decision tree with max_depth up to 4.
  • Bigger – gradient boosted trees with max_depth up to 4 and number of trees limited to 40.
In each case, the data set was split 80:20 into training and test sets. In addition, for each model and its constraint, a grid search was performed over selected hyperparameters. The details of the training process can be found in the Appendix.
Measurements
The comparison was carried out in different settings. For a selected dataset and sampling method (MC/QMC), the variable parameters were:
  • Model type $m$. Recall that, in our case, there were two model types for a fixed dataset.
  • The standard deviation $\sigma$ of a Gaussian used to perturb a feature. We used the values $\{0.1,0.3,1.0\}$.
[…]
Implementation
All algorithms were implemented in Python with the key parts including inference on the tree ensemble being implemented as a C++ package and all the arithmetic operations in the algorithms were carried out using the numpy.float32 type or C++ float type.